# Wavespeed selection of travelling wave solutions of a two-component reaction-diffusion model of cell invasion

### v1

@@ Abstract
We consider a two-component reaction-diffusion system that has previously been developed to model invasion of cells into a resident cell population. This system is a generalisation of the well-studied Fisher–KPP reaction diffusion equation. By explicitly calculating families of travelling wave solutions to this problem, we observe that a general initial condition with either compact support, or sufficiently large exponential decay in the far field, tends to the travelling wave solution that has the largest possible decay at its front. Initial conditions with sufficiently slow exponential decay tend to those travelling wave solutions that have the same exponential decay as their initial conditions. We also show that in the limit that the (nondimensional) resident cell death rate is large, the system has similar asymptotic structure as the Fisher–KPP model with small cut-off factor, with the same universal (leading order) logarithmic dependence on the large parameter. The asymptotic analysis in this limit explains the formation of an interstitial gap (a region preceding the invasion front in which both cell populations are small), the width of which is also logarithmically large in the cell death rate.
@@ 1 Introduction
$\frac{\partial u}{\partial t}=\frac{\partial^{2}u}{\partial x^{2}}+u(1-u).$ (1)
@@
$c=\begin{cases}2&a\geq 1\\ \displaystyle a+\frac{1}{a}&a<1.\end{cases}$
@@
(see also Hamel et al. [16], Nolen et al. [19]). As well as of inherit interest, this slow convergence is important to take into account when trying to estimate the long-time wave speed from numerical simulations.
@@
where $\epsilon\ll 1$. The motivation behind this cut-off term comes from considering (1) as the continuous limit of a discrete process, where $\epsilon$ represents the discrete cell size. The matched asymptotic analysis in [6] established that the correction to the wave speed in the limit $\epsilon\to 0$ was logarithmic:
$c\sim 2-\frac{\pi^{2}}{(\log\epsilon)^{2}},\qquad\epsilon\to 0,$ (4)
@@
To model more complex biological processes such as tumour growth, in which multiple cell or chemical species may be involved, systems of multiple reaction-diffusion models have been developed and analysed. Gatenby and Gawlinski [15] developed a three-species model of tumour growth under the acid-mediation hypothesis, in which an invading cell population increases the acidity of the environment (modelled by the production and diffusion of H+ ions), which then negatively impacts the resident (non-diffusive) healthy cell population [21, 3]. In this model, the diffusivity of the invading cells is assumed to be a decreasing function of the resident cell population, such that when the healthy cell population is at capacity, the diffusivity vanishes. The decrease in resident cell population due to acidity is thus required to allow the tumour to grow.
@@
In Gatenby and Gawlinski [15] it was numerically observed that solutions to their three-component model tend to travelling wave solutions as time increases. A particular point of interest was that when the death of resident cells due to acidity was large, there was a separation between the advancing front of invading cells and receding front of resident cells, in which both cell populations were small. This region is referred to as the hypocellular interstitial gap, and was observed in experimental observations. Gatenby and Gawlinski [15] estimated that the size of this region was logarithmically large in the dimensionless resident cell death rate. The asymptotic analysis of travelling wave solutions has also been carried out more recently in Fasano et al. [12], Davis et al. [8]; the interstitial gap is a feature of waves that have speed $O(1)$ in the limit that acid diffusivity is much larger than the cell diffusivity (referred to as ‘slow’ waves in Fasano et al. [12], Davis et al. [8], the ‘fast’ waves relating to the finite diffusivity of the acid species).
Simplified versions of the three-species model of Gatenby and Gawlinski [15] have been studied, by neglecting the acid species, and directly modelling the effect of resident cell population on the diffusivity of the invading species. Gallay and Mascia [14] show that travelling wave solutions exist for any wavespeed $c$, while Mascia et al. [18] describe numerical solution methods. The stability of travelling wave solutions was considered in Swartwood [22], who show that all travelling waves are stable, regardless of the speed $c$ (this is similar to the Fisher–KPP equation (1), where the marginal stability threshold is considered as the relevant wavespeed selection mechanism [23, 24]). Colson et al. [7] consider the case where in addition, resident cell species do not recover, and may initially be at less than the carrying capacity. In that study, families of travelling wave solutions are constructed in phase space by a shooting algorithm. A minimum wave speed is identified as a function of the resident cell death rate, and the initial resident cell population.
@@
Recently, Browning et al. [5] and El-Hachem et al. [11] consider another version of a simplified model of tumour growth whereby the two cell species react directly. In this model, both the diffusion and the carrying capacity of the invading cells is inhibited by the healthy cell population, and resident cells $v$ do not proliferate, but are only destroyed by the invading species:
@@
In this model, $u$ is the invading species, and $v$ is the resident species. This model may be seen as a version of the system studied in Colson et al. [7] with an additional effect of $v$ on the carrying capacity of $u$. Nominally this model has a single parameter, the dimensionless resident cell death rate $\gamma$, but since in this model there is no proliferation of $v$, the initial condition of $v$ plays an important role. The diffusivity of $u$ vanishes at $v=1$, but the diffusion will only be degenerate if the initial condition takes this value. Although travelling wave solutions of (5) were observed in PDE simulations in El-Hachem et al. [11] over a range of parameter values and initial conditions, and the boundary value problem for travelling wave solutions was posed, the explicit computation of travelling wave solutions was not performed, and the mechanism by which the wave speed is determined by the parameters was not established.
@@
In this article we consider the model (5). We explicitly calculate travelling wave solutions via numerical continuation over a large range of parameter values (in particular the resident cell death rate $\gamma$), and compare with time-dependent numerical simulations. We are thus able to numerically observe that, although the details of the phase space are more complicated, the principle by which the wave speed is selected is the same as for the simpler Fisher–KPP system: that is, an initial condition with sufficiently small exponential decay rate will tend to the travelling wave solution with the same decay rate (if it exists), and that an initial condition with compact support (or sufficiently large exponential decay) will tend to the travelling wave solution that has maximum decay rate. This is similar to the observation made in Colson et al. [7] for their simpler system.
@@
In addition, our method of calculating travelling wave solutions allows us to compute solutions in the asymptotic limit as the resident cell death rate $\gamma$ becomes large. We show that the asymptotic structure in this limit is very similar to that for the Fisher–KPP system with cut off (3), in which $\gamma^{-1}$ plays the role of the small parameter $\epsilon$. The leading order correction to the velocity is the same as the universal correction (4). We also find the next term in the correction, although this requires the numerical calculation of prefactors in each of the leading order problems in the asymptotic analysis. By carrying out a matched asymptotic expansion in this limit, we observe that the wave speed converges logarithmically to the (nondimensional) Fisher–KPP wave speed of $c=2$. This asymptotic solution also explains why the model exhibits an interstitial gap where both cell populations are small; the width of this zone is also logarithmically dependent on the cell death parameter. We show that the predictions made by the asymptotic analysis agree with the numerically computed travelling wave solutions.
@@ 2 Model and numerical PDE simulations
We start by describing the method and results of numerical simulations of the system (5). El-Hachem et al. [11] report on a large number of numerical simulations of this system; our aim here is to reproduce their main observations in order to confirm the validity of the travelling wave solutions in the next section. The system (5) comprises a population density of invading cells $u(x,t)$ invading a stationary population of resident cells $v(x,t)$, where both the diffusion and carrying capacity of the invading cells is negatively affected by the resident population. The nondimensional parameter $\gamma$ in (5) represents the rate of destruction of resident cells, compared to the proliferation time scale; Browning et al. [5] and El-Hachem et al. [11] contain further details of this model, including the nondimensionalisation.
@@
Note that if $v_{0}=0$ identically, the system (5) reduces to the Fisher–KPP problem for $u$. In addition, the diffusivity vanishes when $v=1$; in this paper, we will consider only $10<v_{0}<$ 1, as our main focus is on the effects of $v$ on the proliferation, rather than the effect of degenerate diffusion (see the discussion in Sect. 5).
[…]
@@
where $c$ is the wave speed, and $k_{0}$ and $k_{1}$ are constants; $k_{1}$ depends on the initial condition, and while $k_{0}=-3/2$ for the minimum-speed Fisher–KPP solution, it likely depends on the parameters for the more complex model (5). To estimate $c$ we thus fit the nonlinear relationship (7) for large times. We find that including this logarithmic correction improves the estimate of the long-time wave speed.
In Fig. 1 we plot typical results (profiles of $u$ and $v$ over time) for a combination of different cell death rates $\gamma$ and initial conditions. In Fig. 1(a,b), we consider an initially compactly supported invading cell population:
@@
for the values $\gamma=1$ and $\gamma=10$, respectively. It is apparent from examining the behaviour of solutions that the system is tending toward a travelling wave of fixed shape and constant speed. Using the method described above, we estimate the wave speed of each of these cases to be $c=1.00$ and $c=1.24$, respectively. In Figs. 1(c,d) we consider initial conditions with exponential decay, and larger initial resident cell population:
@@
for $a=0.27$ and $a=0.21$, respectively, and again for $\gamma=1$ and $\gamma=10$, respectively. These decay rates $a$ have been chosen so that 1(c) has the same long-time wave speed as 1(a), and 1(d) has the same long-time wave speed as 1(b). A linear analysis near the front, as carried out in El-Hachem et al. [11], and also below in Sect. 3, predicts that for an exponentially decaying initial condition $u_{0}(x)=\exp(-ax)$, the wave speed is given by
$c=\begin{cases}2(1-v_{0})&a\geq 1\\ \left(a+\frac{1}{a}\right)(1-v_{0})&a<1,\end{cases}$ (8)
which motivates the choice of initial conditions above (we will see, however, that (8) only holds for sufficiently small $\gamma$). What is important to note in Fig. 1 is that the wave speed is dependent on both the parameter $\gamma$ and the initial conditions of $u$ and $v$, and, indeed, two different sets of initial conditions for the same value of $\gamma$ can result in the same wave speed $c$.
@@
To further explore the possible behaviour of the system (5), we run simulations and estimate travelling wave speeds for a much wider range of parameter values and initial conditions. For compactly supported $u_{0}(x)$, we estimate the wave speed for values of $\gamma$ ranging from $0.1$ to $10^{6}$, and for values $v_{0}\in\{0.25,0.5,0.75\}$. These estimated wave speeds are plotted in Fig. 2, and will be used to validate the predictions made from the direct computation of travelling wave solutions in the next section. The most notable property of these speeds is that for sufficiently small $\gamma$, the wave speed is given by $c=2(1-v_{0})$ independent of $\gamma$, while for sufficiently large $\gamma$, the wave speed becomes dependent on $\gamma$. The point of transition between these behaviours depends on $v_{0}$, but in all cases, as $\gamma\to\infty$, the speed $c$ approaches $2$ from below.
@@
Figure 2: The dependence of wave speed $c$ on death rate $\gamma$, for different values of initial resident cell population $v_{0}=V_{\infty}$. Blue symbols indicate the estimation of wave speed from PDE simulations described in Sect. 2. The dashed curves represent the predicted wave speed according to the condition that $V_{\infty}=V_{c}$ (12), while the solid curves represent the predicted wave speed according to the condition $V_{\infty}=V_{s}$, where $V_{s}$ must be found by numerically solving the system (9). The PDE simulations with initial condition $v_{0}=V_{\infty}$ predict the same wave speed as the travelling wave analysis.
@@
We test the effect of initial condition of $u$ by calculating the wave speed for $\gamma\in[0.1,10^{6}]$, fixed $v_{0}=0.5$, and exponentially decaying initial conditions $u_{0}=\mathrm{e}^{-ax}$ for $a\in\{0.5,0.325,0.25\}$. These wave speeds are plotted in Fig. 3. These results show that for sufficiently small $a$ and $\gamma$, the wave speed is given by the formula (8). Furthermore, for a fixed $a$ sufficiently small such that the wave speed $c$ is greater than 2, that is
$a>a^{*}=\frac{1}{1-v_{0}}-\sqrt{\frac{1}{(1-v_{0})^{2}}-1},$
the wave speed is given by (8) for any $\gamma$. The value $a=0.25$ shown in Fig. 3 satisfies this condition. On the other hand, if $a<a^{*}$, for sufficiently large $\gamma$ the wave speed is the same as the $\gamma$-dependent wave speed of the compactly supported initial condition. Equivalently, for a fixed $\gamma$, as $a$ is increased, the wave speed is ultimately limited by the corresponding wave speed of the compactly supported initial condition of the same value of $\gamma$. In the next section we explain this phenomenon via the calculation of travelling wave solutions in phase space.
@@
Figure 3: The dependence of wave speed $c$ on death rate $\gamma$, for initial resident cell population $v_{0}=0.5$ and different invading cell initial conditions $u_{0}=\exp(-ax)$, for $a\in\{0.5,0.325,0.25\}$. The wave speed predicted by the formula (8) is depicted by the coloured dashed lines. The wave speed of a compactly supported initial condition, corresponding to that in Fig. 2, is shown as black dashed/solid lines.

### v2

# Wavespeed selection and interstitial gap formation in an acid-mediated cancer invasion model

@@ Abstract
We consider a two-component reaction-diffusion system that has previously been developed to model invasion of cells into a resident cell population. The system is an idealised version of models of tumour growth in which tumour cells degrade the surrounding tissue by increasing the acidity of the local environment. By numerically computing families of travelling wave solutions to this problem, we observe that a general initial condition with either compact support, or sufficiently large exponential decay in the far field, tends to the travelling wave solution that has the largest possible decay at its front. Initial conditions with sufficiently slow exponential decay tend to those travelling wave solutions that have the same exponential decay as their initial conditions. We also show that in the limit that the (nondimensional) degradation rate of resident cells is large, the system has similar asymptotic structure as previously observed in perturbed Fisher–KPP models. The asymptotic analysis in this limit explains the formation of an interstitial gap (a region between the invading and receding fronts, in which both cell populations are small), the width of which is logarithmically large in the limit of large degradation rate. These results show that the general mechanism behind the formation of the interstitial gap in reaction-diffusion tumour models is connected to perturbations of the Fisher-KPP system. Biologically, this implies that order of magnitude difference in degradation rate is required to produce appreciably different gap sizes, while the velocity of the invading front is largely determined by the Fisher-KPP velocity, and only very weakly affected by the presence of the interstitial gap.
keywords:
Reaction–Diffusion, Travelling waves, tumour growth
@@ 1 Introduction
$\frac{\partial u}{\partial t}=\frac{\partial^{2}u}{\partial x^{2}}+u(1-u),\qquad t>0.$ (1)
Generally speaking, $u$ could model a population of cells, chemical concentration, or the prevalence of a gene, for example, and the problem is considered on an infinite spatial domain ($\infty<x<\infty$), or a semi-infinite domain ($0<x<\infty$) with suitable (usually no-flux) boundary condition.
@@
$c=\begin{cases}2,&a\geq 1,\\ \displaystyle a+\frac{1}{a},&a<1.\end{cases}$
@@
(see also Hamel et al. [17], Nolen et al. [25]). As well as of inherent interest, the knowledge that convergence can be of this form is useful for estimating the long-time wave speed from numerical simulations.
@@
where $0<\epsilon\ll 1$. The motivation behind this cut-off term comes from considering (1) as the continuum limit of a discrete process, where $\epsilon$ represents the discrete cell size. The matched asymptotic analysis in [3] established that the correction to the minimum wave speed (that is, the speed attained by initial conditions with sufficiently fast decay) in the limit $\epsilon\to 0^{+}$ was logarithmic:
$c\sim 2-\frac{\pi^{2}}{(\log\epsilon)^{2}},\qquad\epsilon\to 0^{+},$ (4)
@@
The growth of malignant tumours has long been connected with the tendency of tumour cells to acidify their local extra-cellular environment [5, 27]. This tendency is due to the Warburg effect, by which cancer cells preferentially obtain energy by lactic fermentation, even when oxygen is available. The resulting environment is unfavourable to healthy cells and leads to degradation of the extracellular matrix, resulting in more space for invading cancer cells to invade.
Multiple approaches have been used to mathematically model the process of acid-mediated tumour growth. Most relevant for this study is the use of reaction-diffusion partial differential equation (PDE) systems, as originated in Gatenby and Gawlinski [15]. These are generally macroscopic models in which cell populations and extracellular chemical concentrations are treated as continuous functions of space, and heterogeneities due to tissue structure or vascularity are neglected. The modelling approach of Gatenby and Gawlinski [15] has been subsequently extended in a number of ways, for example to include direct interactions between cancer and healthy cells (that is, separating the dynamics of healthy cells from degradation of extracellular matrix) [22, 23], or the effects of chemotaxis (repulsion of cells from areas of high acidity) [12]. Multiscale models that include the coupling of extra- and intracellular acidity, often with stochastic effects included, have been developed, simulated and long-time existence proven [19, 18, 20].
The reaction-diffusion PDE model of Gatenby and Gawlinski [15] may be written in nondimensional form as
$\frac{\partial u}{\partial t}$ $=\frac{\partial}{\partial x}\left[(1-v)\frac{\partial u}{\partial x}\right]+u(1-u),$ (5a)
$\frac{\partial v}{\partial t}$ $=\gamma_{1}v(1-v-\alpha h),$ (5b)
$D\frac{\partial h}{\partial t}$ $=\gamma_{2}(u-h)+\frac{\partial^{2}h}{\partial x^{2}},$ (5c)
where $u(x,t)$ is the invading (cancer) cell density, $v(x,t)$ the resident (healthy) cell density, and $h(x,t)$ the acid concentration, while $\alpha$, $\gamma_{1}$, $\gamma_{2}$, and $D$ are nondimensional parameters. In this model the resident cell population $v$, which effectively represents the combination of healthy cells and extracellular matrix, is stationary (non-diffusive) and proliferates with a carrying capacity that decreases with the acid concentration. The acid $h$ is produced by the invading cells, while undergoing diffusion and degradation. The parameter $D$ represents the ratio of cell to acid diffusivities, and would generally be taken to be small. Importantly, the diffusivity of the invading cell population is a decreasing function of the resident cell population, such that when the healthy cell population is at capacity, the diffusivity vanishes. The decrease in resident cell population due to acidity is thus required to allow the tumour to grow.
@@
In Gatenby and Gawlinski [15] it was numerically observed that solutions to their three-component model (5) tend to travelling wave solutions as time increases. In such solutions, there is an advancing front of the invading cells $u$ similar to the Fisher–KPP model coupled to a receding front of the resident cells $v$. A particular point of interest was that when the death of resident cells due to acidity is large (large $\alpha$ in (5)), there is a separation between the advancing and receding fronts, in between which both cell populations are small. Such a region corresponds to a phenomenon sometimes observed in biological experiments referred to as the hypocellular interstitial gap. Gatenby and Gawlinski [15] estimated that the size of this region was logarithmically large in $\alpha$. The asymptotic analysis of travelling wave solutions has also been carried out more recently in Fasano et al. [13] and Davis et al. [10]; the interstitial gap is a feature of waves that have speed $O(1)$ in the limit that acid diffusivity is much larger than the cell diffusivity. These waves are referred to as ‘slow’ waves in Fasano et al. [13] Davis et al. [10], the ‘fast’ waves relating to the finite diffusivity ($D>0$ in (5)) of the acid species.
Simplified versions of the three-species model of Gatenby and Gawlinski [15] have been studied, by neglecting the acid species, and directly modelling the effect of resident cell population on the diffusivity of the invading species. Gallay and Mascia [16] and Mascia et al. [24] consider existence of travelling wave solutions for the model
$\frac{\partial u}{\partial t}=\frac{\partial}{\partial x}\left[(1-v)\frac{\partial u}{\partial x}\right]+u(1-u),\qquad\frac{\partial v}{\partial t}=\gamma_{1}v(1-v-\alpha u).$ (6)
Formally this model is recovered from (5) by taking the coupling between invading cell population $u$ and acid concentration $h$ to be strong, that is, $\gamma_{2}\gg 1$, so that $u\sim h$. The stability of travelling wave solutions of this system is considered in Swartwood [28], who show that all travelling waves are stable, regardless of the speed $c$; this is similar to the Fisher–KPP equation (1), where the marginal stability threshold is considered as the relevant wavespeed selection mechanism [29, 30].
Colson et al. [7] consider a version of the reduced model (6) in which the resident cell species does not recover:
$\frac{\partial u}{\partial t}=\frac{\partial}{\partial x}\left[(1-v)\frac{\partial u}{\partial x}\right]+u(1-u),\qquad\frac{\partial v}{\partial t}=-\gamma uv.$ (7)
This may be considered the large-$\alpha$ approximation of (6), identifying $\gamma=\alpha\gamma_{1}$. In this model the initial resident cell population plays an important role, as the initial population may or may not be taken to be the value at which the diffusivity of the invading cell population vanishes ($v=1$, nondimensionally). In Colson et al. [7] families of travelling wave solutions are constructed in phase space by a shooting algorithm. A minimum wave speed is identified as a function of the resident cell death rate, and the initial resident cell population.
@@
Recently, Browning et al. [4] and El-Hachem et al. [11] consider a similar model to (7), in which both the diffusion and the carrying capacity of the invading cells is inhibited by the healthy cell population:
@@
with $t>0$. In the context of travelling waves this model is posed on an infinite domain $\infty<x<\infty$, or equivalently a semi-infinite domain $0<x<\infty$ with appropriate no-flux condition at $x=0$. This model may be seen as a version of the system (7) with an additional effect of resident cell density $v$ on the carrying capacity of invading cells $u$. Nominally this model has a single parameter, the dimensionless resident cell death rate $\gamma$, but since in this model there is no proliferation of $v$, the initial condition of $v$ again plays an important role. The diffusivity of $u$ vanishes at $v=1$, but the diffusion will only be degenerate if the initial condition takes this value. Although travelling wave solutions of (8) were observed in PDE simulations in El-Hachem et al. [11] over a range of parameter values and initial conditions, and the boundary value problem for travelling wave solutions was posed, the explicit computation of travelling wave solutions was not performed, and the mechanism by which the wave speed is determined by the parameters was not established.
@@
In this article we focus on the model (8), with the aim to build insight into not only the properties of travelling wave solutions, but also the selection of a given wave speed from a given initial condition. While this was only explored using PDE simulations in El-Hachem et al. [11], we explicitly calculate travelling wave solutions via numerical continuation over a large range of parameter values (in particular the resident cell death rate $\gamma$), and compare with time-dependent numerical simulations. We are thus able to observe that, although the details of the phase space are more complicated, the principle by which the wave speed is selected is the same as for the simpler Fisher–KPP system: that is, an initial condition with sufficiently small exponential decay rate will tend to the travelling wave solution with the same decay rate (if it exists), and that an initial condition with compact support (or sufficiently large exponential decay) will tend to the travelling wave solution that has maximum decay rate. This is similar to the observation made in Colson et al. [7] for their simpler system (7).
In addition, we explore the asymptotic limit as the resident cell death rate $\gamma$ becomes large, which was not achieved in El-Hachem et al. [11]. As discussed above, it is in this limit that we may connect the hierarchy of models (8) and (7) to (6) and ultimately the complete Gatenby–Gawlinski model (5). In addition, this limit corresponds to the large-$\alpha$ limit of the Gatenby–Gawlinski model in which we expect to see a separation of advancing and receding fronts. Importantly, in this limit each of the above models becomes a perturbed version of the Fisher–KPP equation (1), in that the equation for invading cells $u$ reduces to (1) when the resident population $v$ vanishes.
@@
We demonstrate that there is a very strong connection between the asymptotic limit of large resident cell death rate and the analysis of the Fisher–KPP system with cut off (3) [3, 9], with $\gamma^{-1}$ playing the role of the small parameter $\epsilon$. The leading order correction to the velocity is the same as the universal correction (4). We also find the next term in this relation, although this requires the numerical calculation of prefactors in each of the leading order problems in the asymptotic analysis. This asymptotic solution also exhibits a region between fronts where both populations are small, that is, an interstitial gap; the width of this zone is also logarithmically dependent on the cell death parameter. As our numerical continuation method allows us to numerically compute travelling wave solutions up to very large ($10^{11}$) values of $\gamma$, we are able to numerically observe the correctness of the asymptotic approximation. The universality of this asymptotic approach demonstrates the fundamental dynamics that is at play in interstitial gap formation in the large family of reaction-diffusion models of acid-mediated cancer invasion.
@@ 2 Model and numerical PDE simulations
We start by describing the method and results of numerical simulations of the system (8). El-Hachem et al. [11] report on a large number of numerical simulations of this system; our aim in this section is to reproduce their main observations in order to confirm the validity of the travelling wave solutions in the next section. The system (8) comprises a population density of invading cells $u(x,t)$ invading a stationary population of resident cells $v(x,t)$, where both the diffusive flux and carrying capacity of the invading cells is negatively affected by the resident population. The nondimensional parameter $\gamma$ in (8) represents the rate of destruction of resident cells, compared to the time scale of invading cell proliferation. Browning et al. [4] and El-Hachem et al. [11] contain further details of this model, including the nondimensionalisation.
@@
Note that if $v_{0}=0$ identically, the system (8) reduces to the Fisher–KPP problem for $u$. In addition, the diffusivity vanishes when $v=1$; in this paper, we will consider only $0<v_{0}<1$, as our main focus is on the effects of $v$ on the proliferation, rather than the effect of degenerate diffusion (see the discussion in Sect. 5).
To simulate (8) numerically, we discretise in space using a standard cell-centred finite volume method [26] with equally sized cells, and advance in time using the ode15s algorithm in MATLAB. To observe the formation of travelling wave behaviour, we consider a large spatial domain $x\in[0,L]$ with $L\gg 1$ approximating a semi-infinite domain, with zero flux conditions ($\partial_{x}u=0$) conditions imposed at the boundaries. A nonzero initial condition $u(x,0)$ near $x=0$ is then used to represent a localised initial condition. We generally choose $L$ to be between $150$ and $200$ and $N=10^{4}$ grid points, which is more than adequate for the numerical simulations to have converged.
@@
where $c$ is the wave speed, and $k_{0}$ and $k_{1}$ are constants; $k_{1}$ depends on the initial condition, and while $k_{0}=-3/2$ for the minimum-speed Fisher–KPP solution, it likely depends on the parameters for the more complex model (8). To estimate $c$ we thus fit the nonlinear relationship (10) for large times using a least-squares procedure.
In Fig. 1 we plot typical results (profiles of $u$ and $v$ over time) for a combination of different cell death rates $\gamma$ and initial conditions. We choose specific combinations of $\gamma$, $v_{0}$, and either compactly supported or exponentially decaying $u_{0}(x)$, in order to demonstrate the nontrivial dependence of the wave speed on these parameters that will be further explored using the travelling wave analysis in the next section.
In Fig. 1(a,b), we consider an initially compactly supported invading cell population:
@@
for the values $\gamma=1$ and $\gamma=10$, respectively. It is apparent from examining the behaviour of solutions that the system is tending toward a travelling wave of fixed shape and constant speed. The speed $c$ depends on the value of $\gamma$; applying the fitting procedure described above to the function (10), we estimate $c=1.00$ for $\gamma=1$, and $c=1.24$ for $\gamma=10$. We note that these solutions have wave speed considerably less than the minimum wave speed $c=2$ of the Fisher–KPP system (1), although $c$ is seen to increase as $\gamma$ increases.
In Figs. 1(c,d) we consider initial conditions with exponential decay, and larger initial resident cell population $v_{0}$:
@@
for $a=0.27$ and $a=0.21$, respectively, and again for $\gamma=1$ and $\gamma=10$, respectively. These solutions are seen to converge to travelling waves with the same respective wave speeds: $c=1$, and $c=1.24$, respectively. These results show that different combinations of initial conditions can lead to the same wave speed. We will further explain this relationship in the travelling wave analysis of the next section, using these parameter values as specific examples.
@@
To further demonstrate the possible behaviour of the system (8), we run simulations and estimate travelling wave speeds for a much wider range of parameter values and initial conditions. For compactly supported $u_{0}(x)$, we estimate the wave speed for values of $\gamma$ ranging from $0.1$ to $10^{6}$, and for values $v_{0}\in\{0.25,0.5,0.75\}$. These estimated wave speeds are plotted in Fig. 2, and will be used to validate the predictions made from the direct computation of travelling wave solutions in the next section. The most notable property of these speeds is that for sufficiently small $\gamma$, the wave speed is independent of $\gamma$, while for sufficiently large $\gamma$, the wave speed becomes dependent on $\gamma$. The point of transition between these behaviours depends on $v_{0}$, but in all cases, as $\gamma\to\infty$, the speed $c$ approaches $2$ from below.
@@
Figure 2: The dependence of wave speed $c$ on death rate $\gamma$, for different values of initial resident cell population $v_{0}=V_{\infty}$. Blue symbols indicate the estimation of wave speed from PDE simulations with compactly support initial conditions, described in Sect. 2. The dashed curves represent the predicted wave speed according to the condition that $V_{\infty}=V_{c}$ (16), while the solid curves represent the predicted wave speed according to the condition $V_{\infty}=V_{s}$, where $V_{s}$ must be found by numerically solving the system (11). The PDE simulations with initial condition $v_{0}=V_{\infty}$ predict the same wave speed as the travelling wave analysis.
@@
Finally, we further test the effect of initial condition of $u$ by calculating the wave speed for $\gamma\in[0.1,10^{6}]$, fixed $v_{0}=0.5$, and exponentially decaying initial conditions $u_{0}=\mathrm{e}^{-ax}$ for $a\in\{0.5,0.325,0.25\}$. These wave speeds are plotted in Fig. 3. These results show that for sufficiently small $a$ and $\gamma$, the wave speed is again independent of $\gamma$. Furthermore, for a fixed $a$ sufficiently small such that the wave speed $c$ is greater than 2, for example $a=0.25$ shown in Fig. 3, the wave speed is independent of $\gamma$ over all $\gamma$. On the other hand, if $a$ is sufficiently large such that the speed is less than $2$, for sufficiently large $\gamma$ the wave speed is the same as the $\gamma$-dependent wave speed of the compactly supported initial condition. Equivalently, for a fixed $\gamma$, as $a$ is increased, the wave speed is ultimately limited by the corresponding wave speed of the compactly supported initial condition of the same value of $\gamma$. These phenomena will also be explained by the travelling wave analysis in the next section.
@@
Figure 3: The dependence of wave speed $c$ on death rate $\gamma$, for initial resident cell population $v_{0}=0.5$ and different invading cell initial conditions $u_{0}=\exp(-ax)$, for $a\in\{0.25,0.325,0.5\}$. The symbols represent numerical solutions for $a=0.25$ (filled circles), $a=0.325$ (unfilled circles), and $a=0.5$ (stars), respectively. The wave speed predicted by the formula (17) for each $a$ is depicted by the coloured dashed lines. The wave speed of a compactly supported initial condition, corresponding to that in Fig. 2, is shown as black dashed/solid lines.